integral curve of dx/y^2+yz+x^2= dy/y^2-xz+x^2=dz/z(x+y)
1
Expert's answer
2021-08-27T10:44:34-0400

Solution;

1)Subtract first expression from the second and equate to the third;

"\frac{dx-dy}{y^2+yz+x^2-y^2+xz-x^2}" ="\frac{dz}{z(x+y)}"

To obtain;

"\frac{dx-dy}{z(y+x)}=\frac{dz}{z(x+y)}"

Simplify to obtain;

"d(x-y)=1dz"

Integrate both sides;

"\int d(x-y)=\int 1dz"

"x-y+C=z"

"z=x-y+C_1"

Gives the first integral curve.

2)Add the third expression to the second and equate with the first one;

"\frac{dx}{y^2+yz+x^2}=\frac{dy+dz}{y^2-xz+x^2+xz+yz}"

Simplify;

"\frac{dx}{y^2+yz+x^2}=\frac{dy+dz}{y^2+yz+x^2}"

Multiple both sides with "y^2+yz+x^2" to obtain;

"dx=dy+dz"

"dx=d(y+z)"

Integrate both sides;

"\int dx=\int d(y+z)"

"x+C_2=y+z"

"z=x-y+C_2"

Means the equation has only one integral curve;

"z-x+y=C"